OpenStudy (anonymous):

The volume and surface area of a cylinder are given, respectively, by the formula V=(pie)r^2h and SA=2(pie)h=2(pie)r^2 determine an expression, in simplified form that represents the surface area-to-volume ratio for a cylinder

OpenStudy (nubeer):

ratio = SA/V you ahve SA and V both given just plug it in.

OpenStudy (anonymous):

yes but i couldn't get the right answer

OpenStudy (nubeer):

what is answer

OpenStudy (anonymous):

2h+2r / hr

OpenStudy (anonymous):

your surface area should be SA = \(\large 2\pi r h + \pi r^2 \)